geerky42 (geerky42):

A time-varying net force acting on a 4.7 kg particle causes the object to have a displacement given by x = a + b t + d t²+ e t³, where a = 0.83 m , b = 2.6 m/s, d = −3.4 m/s², and e = 1 m/s³, with x in meters and t in seconds. Find the work done on the particle in the first 2.7 s of motion.

geerky42 (geerky42):

any idea? @gerryliyana

geerky42 (geerky42):

@Algebraic!

OpenStudy (anonymous):

x = a + b t + d t²+ e t³, v =dx/dt = b + 2dt + 3et^2 a = dv/dt = 2d + 6et a = 2(-3.4) + 6 (1) (2.7) a = -6.8 + 16.2 a = 9.4 F = m a F = (4.7) 9.4 F = 44.18 N W = F x W = 44.18 (a + b t + d t²+ e t³)

OpenStudy (anonymous):

W = 44.18 x (0.83 +(2.6 x 2.7) + (-3.4 x 2.7^2) + (1 x 2.7^3)

OpenStudy (anonymous):

W = 44.18 x 2.747 W =121.36 J

OpenStudy (anonymous):

@geerky42 this my idea

OpenStudy (anonymous):

dx = (b+ 2dt +3et^2) dt F = ma = m*(2d+6et) integrate: m*(2d+6et) *(b+ 2dt +3et^2) dt from t=0..2.7

geerky42 (geerky42):

i don't see what's wrong with your work, but your answer is incorrect.

geerky42 (geerky42):

thanks, @gerryliyana and @Algebraic!

OpenStudy (anonymous):

try the count returned by you in the same way welcome

OpenStudy (anonymous):

Maybe I made a mistake in the calculation

geerky42 (geerky42):

@Algebraic! integrate: " m*(2d+6et) *(b+ 2dt +3et^2) dt" why (b+ 2dt +3et^2)? I though we only have to integrate the force itself? F = m(2ct + 3et²), right?

OpenStudy (anonymous):

F*dx

OpenStudy (anonymous):

and F is ma ... what's a?

OpenStudy (anonymous):

not (2ct +3et^2) ...

geerky42 (geerky42):

well, W = Fx i have to integrate x too? it is 2c + 6et m(2ct +3et^2) is integral of force, wrong expression, my bad

OpenStudy (anonymous):

no. W = integral(F*dx)

geerky42 (geerky42):

(assume c = d) I wrote the different way. sorry for confusion -_-

OpenStudy (anonymous):

you're not given F in terms of x though...

OpenStudy (anonymous):

so you have to 'fix' that... by expressing everything in terms of t

OpenStudy (anonymous):

(2d+6et) is acceleration

geerky42 (geerky42):

so F = m(2d + 6et)

OpenStudy (anonymous):

yes

geerky42 (geerky42):

so i have to integrate force in respect to x? if no, why dx?

OpenStudy (anonymous):

'normally' you do... but you can't here... at least not directly. but you can write F(x) in terms of t, and dx in terms of t...

OpenStudy (anonymous):

and then integrate that over the time interval

OpenStudy (anonymous):

Is this going over your head? what class is this problem from?

OpenStudy (anonymous):

calc. based physics I hope..?

geerky42 (geerky42):

ohh, i see. i got it now. thanks.

OpenStudy (anonymous):

sure:) hope it helped..

geerky42 (geerky42):

it's clear and it does help me again, thanks.

OpenStudy (anonymous):

great!